# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 3.1 Finding a Relevant Direction
From the rotational invariance of the Gaussian without loss of generality, we can denote by $\mathbf{e}_{1},\dots,\mathbf{e}_{k^{\prime}}$ an orthonormal basis of the space of $U$ projected onto $E_{S}^{\perp}$. We have that $k^{\prime}\geq 1$ otherwise we would not have the existence of $q_{S}$ that is defined over this space. Furthermore, since $U$ is a subset of $V^{\perp}$ we similarly have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)]>\sigma$ for some $z\in\mathcal{Y}$. Decomposing $q_{S}(\mathbf{x}^{U})$ to the basis of Hermite polynomials we have that $q_{S}(\mathbf{x}^{U})=\sum_{\beta\in\mathbb{N}^{d^{\prime}},1\leq\left\|\beta\right\|_{1}\leq m}\hat{q}_{S}(\beta)H_{\beta}(\mathbf{x}^{U})$ where $d^{\prime}=\dim(U)$. Moreover, note that since $q_{S}$ has no component outside $\mathbf{e}_{1},\dots,\mathbf{e}_{k^{\prime}}$ we have that $q_{S}(\mathbf{x}^{U})=\sum_{\beta\in J}\hat{q}_{S}(\beta)H_{\beta}(\mathbf{x}^{U})$, where $J$ denotes the set of $\beta\in\mathbb{N}^{d^{\prime}}$ such that $\beta_{i}\geq 1$ for some $i\in[k^{\prime}]$. Considering the correlation of $q_{S}$ with the label region $z$ we have that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)]$ $=\sum_{\beta\in J}\hat{q}_{S}(\beta)\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]$
$\leq\left(\sum_{\beta\in J}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]^{2}\right)^{1/2}\;,$
@@
where we used the fact that $\|q_{S}\|_{2}=1$ and the Cauchy-Schwarz inequality. Hence, we have that the sum of the squares of the Hermite coefficients for of the degree $m$ restriction the random variable $\mathds{1}(y=z)$ is at least $\sigma^{2}$.
@@
$\sum_{i=1}^{k^{\prime}}\mathbf{e}_{i}^{\top}\mathbf{M}\mathbf{e}_{i}$ $=\sum_{i=1}^{k^{\prime}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\nabla p_{S,z}(\mathbf{x})\cdot\mathbf{e}_{i})^{2}]\geq\sum_{\beta\in\mathbb{N}^{d}}(\sum_{i=1}^{k^{\prime}}\beta_{i})(\hat{p}_{S,z}(\beta))^{2}$
$\geq\sum_{\beta\in J}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]^{2}-2\eta$
@@
Since $\mathbf{M}$ is symmetric and therefore diagonalizable, we can denote by $\mathbf{u}^{(i)}$ the unit eigenvectors of $\mathbf{M}$ and by $\lambda_{i}$ the corresponding eigenvalues. Additionally, let $U_{S,z}$ denote the set of eigenvectors of $\mathbf{M}$ with eigenvalues greater than ${\sigma^{2}}/{4K}$ (see Line 5 of Algorithm 4). Decomposing $\mathbf{e}_{j}$ to the basis of eigenvectors gives us
$\frac{\sigma^{2}}{4K}+\sum_{\mathbf{u}^{(i)}\in U_{S,z}}\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sum_{i=1}^{d-k}\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\frac{\sigma^{2}}{2K}\;.$
@@
Hence, there exists an eigenvector $\mathbf{u}^{(i)}\in U_{S,z}$ such that $\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sigma^{2}/(4K\lvert U_{S,z}\rvert)$ thus $(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sigma^{2}/(4\lvert U_{S,z}\rvert K\text{tr}(\mathbf{M}))$ as $\mathbf{M}$ is positive definite. Therefore, by B.3 we have that $\lvert U_{S,z}\rvert=\mathrm{poly}(mK/\sigma),\text{tr}(\mathbf{M})=O(m)$ which implies that $\lvert\mathbf{e}_{j}\cdot\mathbf{u}^{(i)}\rvert\geq\mathrm{poly}(\sigma/(mK))$. Finally, note that since $\mathbf{e}_{j}\cdot\mathbf{u}^{(i)}\neq 0$, we have that $\mathbf{u}^{(i)}$ cannot belong in the space $E_{S}$. Thus, there exists a unit vector $\mathbf{w}\in W$ such that $\lvert\mathbf{u}^{(i)}\cdot\mathbf{w}\rvert\geq\epsilon/(\sqrt{\lvert U_{S}\rvert}\Gamma)\geq\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)$, which concludes the proof of Claim 3.9. ∎
@@
First note that by the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]{=(\epsilon^{\prime}/k)^{\Omega(k)}}$ and $\lvert\mathcal{S}\rvert{=(k/\epsilon^{\prime})^{O(k)}}$. Therefore, by union bound and Hoeffding’s inequality, it holds that if $N\geq{(k/\epsilon^{\prime})^{Ck}}\log(\lvert\mathcal{S}\rvert/\delta)={(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$. Hence, it is true that $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N{(\epsilon^{\prime}/k)^{\Omega(k)}}$.
@@
Therefore, for $N\geq(dm)^{Cm}{(k/\epsilon^{\prime})^{Ck}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}}$ for a sufficiently large universal constant $C>0$, by applying Claim 3.9, we have that for all $S\in\mathcal{T}$ it holds that $\mathbf{v}\cdot(\mathbf{w}^{(i)})^{\perp V}\geq{\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)}$ for some $\mathbf{v}\in U_{S},i\in[K]$. Hence, we have that there exists a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ and $i\in[K]$ with $\sum_{S\in\mathcal{T^{\prime}}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\sum_{S\in\mathcal{T^{\prime}}}\operatorname*{\mathbf{Pr}}_{D}[S]/2=\Omega(\alpha/K)$ such that for all $S\in\mathcal{T^{\prime}}$ we have that $\mathbf{v}\cdot(\mathbf{w}^{(i)})^{\perp V}\geq{\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)}$ for some $\mathbf{v}\in U_{S}$.
@@
Denote by $\widehat{\mathbf{U}}$ the matrix computed at Line 6 of Algorithm 4. For some $i\in[K]$ we have that
@@
where we used the fact that $\widehat{\mathbf{U}}$ is psd.
@@
In this section, given Proposition 3.7, we proceed to the proof of Theorem 3.3. Recall that, in Proposition 3.7, we have shown that if our current approximation to the hidden subspace is not accurate enough to produce a classifier that has sufficiently small error, then Algorithm 4 efficiently finds a direction that has non-trivial correlation to the hidden subspace. In the proof that follows, we iteratively apply this argument to show that, after a moderate number of iterations, Algorithm 3 outputs a classifier with sufficiently small error.
@@
Denote by $f^{*}\in\mathcal{F}$ a classifier such that $\operatorname*{\mathbf{Pr}}[f^{*}(\mathbf{x})\neq y]\leq\mathrm{OPT}+\epsilon/4$. We show that Algorithm 3, with high probability, returns a hypothesis $h$ with 0-1 error at most $\tau+\mathrm{OPT}+\epsilon$. Let $W^{*}$ be a subspace of dimension $K$ such that $f^{*}(\mathbf{x})=f^{*}(\mathbf{x}^{W^{*}})$ and denote by $\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)}\in\mathbb{R}^{d}$ an orthonormal basis of $W^{*}$. Let $L_{t}$ be the list of vectors updated by the algorithm (Line 3c of Algorithm 3) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$. Also, let $\epsilon^{\prime}$ be the partition width parameter set at Line 2 of Algorithm 3, and for $t\in[T]$ let $\mathcal{S}_{t}$ be arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2). Let $h_{t}:\mathbb{R}^{d}\to[K]$ be a piecewise constant classifiers, defined as $h_{t}=h_{\mathcal{S}_{t}}$ according to Definition 2.3 for the distribution $D$.
@@
Note that from Lines 1 and 3 of Algorithm 3, we perform at most $\mathrm{poly}(mK\Gamma{\lvert\mathcal{Y}\rvert}/(\epsilon\alpha\sigma))$ iterations. Furthermore, in each iteration, we update the vector set with at most $\mathrm{poly}(mK\Gamma{\lvert\mathcal{Y}\rvert}/(\epsilon\alpha\sigma))$ vectors. Hence, it follows that $k_{t}\leq\mathrm{poly}(mK\Gamma{\lvert\mathcal{Y}\rvert}/(\epsilon\alpha\sigma))$, for all $t=1,\dots,T$.
Assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon/2$ for all $t=1,\dots,T$. Using the fact that $N={d}^{Cm}2^{(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))^{C}}\log(1/\delta)$ for a sufficiently large universal constant $C>0$ (Line 2 of Algorithm 3), we can apply Proposition 3.7 and conclude that, with probability $1-\delta$, there exists unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ for $t=[T]$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$. Thus, by Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$, $\Phi_{t}\leq\Phi_{t-1}-\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$. After $T$ iterations, it follows that $\Phi_{T}\leq\Phi_{0}-T\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))=K-T\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$. However, since $T$ is set to be a sufficiently large polynomial of $m,\Gamma,K,{\lvert\mathcal{Y}\rvert},1/\epsilon,1/\alpha$ and $1/\sigma$ we would arrive at a contradiction, since $\Phi_{T}\geq 0$. Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon/2$, for some $t\in\{1,\dots,T\}$. Since the error of $h_{t}$ can only be decreasing by Claim 2.17 and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
Sample and Computational Complexity: From the analysis above we have that the algorithm terminates in $\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$ iterations and at each iteration we draw of the order of ${d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$ samples. Hence, we have that the total number of samples is ${d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$. Moreover, we use at most $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time as all operations can be implemented in polynomial time. ∎

### v2

@@ 3.1 Finding a Relevant Direction
Next, by the rotational invariance of the Gaussian distribution, we may, without loss of generality, choose an orthonormal basis $\mathbf{e}_{1},\dots,\mathbf{e}_{k^{\prime}}$ for the subspace $U$ projected onto $E_{S}^{\perp}$. Note that $k^{\prime}\geq 1$, otherwise, the polynomial $q_{S}$ (which is defined over this space) could not exist. Furthermore, since $U$ is a subset of $V^{\perp}$, we similarly have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)]>\sigma$ for some $z\in\mathcal{Y}$. Decomposing $q_{S}(\mathbf{x}^{U})$ in the basis of Hermite polynomials we have that $q_{S}(\mathbf{x}^{U})=\sum_{\beta\in\mathbb{N}^{d^{\prime}},1\leq\left\|\beta\right\|_{1}\leq m}\widehat{q}_{S}(\beta)H_{\beta}(\mathbf{x}^{U})$, where $d^{\prime}=\dim(U)$. Moreover, note that since $q_{S}$ has no component outside $\mathbf{e}_{1},\dots,\mathbf{e}_{k^{\prime}}$, we have that $q_{S}(\mathbf{x}^{U})=\sum_{\beta\in J}\widehat{q}_{S}(\beta)H_{\beta}(\mathbf{x}^{U})$, where $J$ denotes the set of $\beta\in\mathbb{N}^{d^{\prime}}$ such that $\beta_{i}\geq 1$ for some $i\in[k^{\prime}]$. Considering the correlation of $q_{S}$ with the region of the label $z$, we have that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)]$ $=\sum_{\beta\in J}\widehat{q}_{S}(\beta)\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]$
$\leq\left(\sum_{\beta\in J}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]^{2}\right)^{1/2}\;,$
@@
where we used the fact that $\|q_{S}\|_{2}=1$ and the Cauchy-Schwarz inequality. Hence, we deduce that the sum of the squares of the Hermite coefficients corresponding to degrees up to $m$ in the expansion of the random variable $\mathds{1}(y=z)$ is at least $\sigma^{2}$.
@@
$\sum_{i=1}^{k^{\prime}}\mathbf{e}_{i}^{\top}\mathbf{M}\mathbf{e}_{i}$ $=\sum_{i=1}^{k^{\prime}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\nabla p_{S,z}(\mathbf{x})\cdot\mathbf{e}_{i})^{2}]\geq\sum_{\beta\in\mathbb{N}^{d}}\left(\sum_{i=1}^{k^{\prime}}\beta_{i})(\widehat{p}_{S,z}(\beta)\right)^{2}$
$\geq\sum_{\beta\in J}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]^{2}-2\eta$
@@
Since $\mathbf{M}$ is symmetric and therefore diagonalizable, we can denote by $\mathbf{u}^{(i)}$ the unit eigenvectors of $\mathbf{M}$ and by $\lambda_{i}$ the corresponding eigenvalues. Additionally, let $U_{S,z}$ denote the set of eigenvectors of $\mathbf{M}$ with eigenvalues greater than ${\sigma^{2}}/{4K}$ (see Line 5 of Algorithm 4). Decomposing $\mathbf{e}_{j}$ in the basis of eigenvectors gives us
$\mathbf{e}_{j}^{\top}\mathbf{M}\,\mathbf{e}_{j}=\sum^{d-k}_{i=1}\lambda_{i}\,(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}.$
Using that $U_{S,z}$ contains eigenvectors of $\mathbf{M}$ with eigenvalues greater than ${\sigma^{2}}/{4K}$, we get that
$\frac{\sigma^{2}}{4K}+\sum_{\mathbf{u}^{(i)}\in U_{S,z}}\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\mathbf{e}_{j}^{\top}\mathbf{M}\,\mathbf{e}_{j}\geq\frac{\sigma^{2}}{2K}\;.$
@@
Hence, there exists an eigenvector $\mathbf{u}^{(i)}\in U_{S,z}$ such that $\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sigma^{2}/(4K\lvert U_{S,z}\rvert)$. This implies that $(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sigma^{2}/(4\lvert U_{S,z}\rvert K\text{tr}(\mathbf{M}))$ as $\mathbf{M}$ is positive definite. Therefore, by B.3, we have that $\lvert U_{S,z}\rvert=\mathrm{poly}(mK/\sigma),\text{tr}(\mathbf{M})=O(m)$ which implies that $\lvert\mathbf{e}_{j}\cdot\mathbf{u}^{(i)}\rvert\geq\mathrm{poly}(\sigma/(mK))$. Finally, note that since $\mathbf{e}_{j}\cdot\mathbf{u}^{(i)}\neq 0$, we have that $\mathbf{u}^{(i)}$ cannot belong in the space $E_{S}$. Thus, there exists a unit vector $\mathbf{w}\in W$ such that $\lvert\mathbf{u}^{(i)}\cdot\mathbf{w}\rvert\geq\epsilon/(\sqrt{\lvert U_{S}\rvert}\Gamma)\geq\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)$, which concludes the proof of Claim 3.9. ∎
@@
By the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]=(\epsilon^{\prime}/k)^{\Omega(k)}$ and $\lvert\mathcal{S}\rvert=(k/\epsilon^{\prime})^{O(k)}$. Therefore, by the union bound and Hoeffding’s inequality, it holds that if the sample size $N\geq(k/\epsilon^{\prime})^{Ck}\log(\lvert\mathcal{S}\rvert/\delta)=(k/\epsilon^{\prime})^{O(k)}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$. Hence, $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N(\epsilon^{\prime}/k)^{\Omega(k)}$.
@@
If we choose $N$ so that $N\geq(dm)^{Cm}(k/\epsilon^{\prime})^{Ck}\log(1/\delta)(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}$ for a sufficiently large universal constant $C>0$, by Claim 3.9, we have that for all $S\in\mathcal{T}$ it holds that $\mathbf{v}\cdot(\mathbf{w}^{(i)})^{\perp V}\geq\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)$ for some $\mathbf{v}\in U_{S},i\in[K]$. Hence, we have that there exists a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ and $i\in[K]$ with $\sum_{S\in\mathcal{T^{\prime}}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\sum_{S\in\mathcal{T^{\prime}}}\operatorname*{\mathbf{Pr}}_{D}[S]/2=\Omega(\alpha/K)$ such that for all $S\in\mathcal{T^{\prime}}$ we have that $\mathbf{v}\cdot(\mathbf{w}^{(i)})^{\perp V}\geq\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)$ for some $\mathbf{v}\in U_{S}$.
@@
Denote by $\widehat{\mathbf{U}}$ the matrix computed in Line 6 of Algorithm 4. For some $i\in[K]$, we have that
@@
where we used the fact that $\widehat{\mathbf{U}}$ is a PSD matrix.
@@
In this section, given Proposition 3.7, we prove Theorem 3.3. Recall that, in Proposition 3.7, we have shown that if our current approximation to the hidden subspace is not accurate enough to produce a classifier that has sufficiently small error, then Algorithm 4 efficiently finds a direction that has non-trivial correlation to the hidden subspace. In the proof that follows, we iteratively apply this argument to show that, after a moderate number of iterations, Algorithm 3 outputs a classifier with sufficiently small error.
@@
Denote by $f^{*}\in\mathcal{F}$ a classifier such that $\operatorname*{\mathbf{Pr}}[f^{*}(\mathbf{x})\neq y]\leq\mathrm{OPT}+\epsilon/4$. We show that Algorithm 3, with high probability, returns a hypothesis $h$ with 0-1 error at most $\tau+\mathrm{OPT}+\epsilon$. Let $W^{*}$ be a subspace of dimension $K$ such that $f^{*}(\mathbf{x})=f^{*}(\mathbf{x}^{W^{*}})$ and denote by $\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)}\in\mathbb{R}^{d}$ an orthonormal basis of $W^{*}$. Let $L_{t}$ be the list of vectors maintained by the algorithm (Line 3c of Algorithm 3) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$. Let $\epsilon^{\prime}$ be the partition width parameter set at Line 2 of Algorithm 3. For $t\in[T]$, let $\mathcal{S}_{t}$ be arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2). Let $h_{t}:\mathbb{R}^{d}\to[K]$ be piecewise constant classifiers, defined as $h_{t}=h_{\mathcal{S}_{t}}$ according to Definition 2.3 for the distribution $D$.
@@
Note that from Lines 1 and 3 of Algorithm 3, we perform at most $\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$ iterations. Furthermore, in each iteration, we update the vector set with at most $\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$ vectors (see Proposition 2.8). Hence, it follows that $k_{t}\leq\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$, for all $t=1,\dots,T$.
Assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon/2$ for all $t=1,\dots,T$. Using the fact that $N={d}^{Cm}2^{(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))^{C}}\log(1/\delta)$ for a sufficiently large universal constant $C>0$ (Line 2 of Algorithm 3), we can apply Proposition 3.7 and conclude that, with probability $1-\delta$, there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ for $t=[T]$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$. Thus, by Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$,
$\Phi_{t}\leq\Phi_{t-1}-\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))\;.$
After $T$ iterations, it follows that $\Phi_{T}\leq\Phi_{0}-T\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))=K-T\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$. However, since $T$ is set to be a sufficiently large polynomial of $m,\Gamma,K,\lvert\mathcal{Y}\rvert,1/\epsilon,1/\alpha$ and $1/\sigma$ we would arrive at a contradiction, since $\Phi_{T}\geq 0$. Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon/2$, for some $t\in\{1,\dots,T\}$. Since the error of $h_{t}$ can only be decreasing by Claim 2.17 and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
Sample and Computational Complexity: From the analysis above we have that the algorithm terminates in $\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$ iterations and at each iteration we draw of order ${d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$ samples. Hence, we have that the total number of samples is ${d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$. Moreover, we use at most $\mathrm{poly}(N)$ time as all operations can be implemented in sample polynomial time. ∎